OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution.

OpenStudy (anonymous):

Let x = Connie's votes Let y = Carl's votes We know that y + 53 = x We also know y + x = 221 We can substitute x into the bottom equation: y + (y + 53) = 221 2y + 53 = 221 2y = 168 y = 84 Therefore, Carl has 84 votes. Connie has 53 more than that, so 84 + 53 = 137. Connie has 137 votes and Carl has 84 votes.

OpenStudy (anonymous):

Lol is there a way i can give you more than 1 medal? haha just kidding but thank you so much